Systems of Equations: Elimination Method
25x = 91 - 16y
16y = 64 - 16x

Answer by HyperBrain(694)

25x = 91 - 16y
16y = 64 - 16x >> 64-16x=16y
Add the two equations:
25x+64-16x=91
9x+64=91
9x=27
x=3
25(3)=91-16y
75=91-16y
16y=16
y=1
Power up,
HyperBrain!
Click here and here